# RT-GuIDE: Real-Time Gaussian splatting for Information-Driven Exploration

### v2

@@ IV-B1 High-level planner
In contrast to traditional mapping representations, the Gaussian map does not encode occluded and free space. Instead of computing (geometric) frontiers, we use the Gaussian uncertainty estimates to identify regions of the map that should be visited next. We evenly partition the a priori known enclosing space of the desired map into cuboidal regions. At time ${k}$, for a region $o$, denote the cardinality of the Gaussians in the region by $M_{k}$. We compute the mean uncertainty of that region $\Omega_{k}=1/{M_{k}}\left(\sum_{\mu\in o}\|\mu_{k}-\mu_{{k-1}}\|_{2}\right).$ We formulate a high-level guidance path that allows us to (i) navigate to regions of high uncertainty; and (ii) utilize the known traversable space to plan long-range trajectories without computationally expensive collision checks. At the high level of our planner, we construct a tree by incrementally adding nodes along the traveled path. The tree consists of two types of nodes: odometry nodes and viewpoint nodes. At each planning iteration, we sample a fixed number of viewpoints around the identified regions. We compute the shortest viewing distance from the Gaussian region centroid to the optical center of the camera, given the camera intrinsics. Each sampled viewpoint is then assigned the utility computed for that region and connected to the closest odometry node in the tree as a viewpoint node. We then use Dijkstra’s algorithm to find the shortest path from the current robot location in the tree to all the viewpoint nodes in the tree. Finally, we compute the cost-benefit of a path using $\Omega/e^{d}$ where $\Omega$ is the estimated information gain and $d$ is the distance to the node [44]. The maximal cost-benefit path is sent to the trajectory planner.
@@ IV-B2 Trajectory planner
$\xi(x,\hat{G})=\sum_{g\in\hat{G}_{H}}v(x,g)-\lambda_{\xi}\sum_{g\in\hat{G}_{L}}v(x,g).$ (9)
The first term in eq. (9) encourages additional observations of parts of the scene with high uncertainty. The second term penalizes observation of stable Gaussians in explored areas and encourages exploration of unseen parts of the environment. The two terms are weighted by the $\lambda_{\xi}$.
@@
Trajectory Candidate Generation. Given an initial robot state $x_{0}\in\mathcal{X}_{free}$, and goal region $\mathcal{X}_{goal}$, find the control inputs $u(\cdot)$ defined on $[0,\tau]$ that solve:
[…]
@@
Motion Primitive Tree Generation. Inspired by [45], we solve problem 1 by performing a tree search on the motion primitives tree. Motion primitives are generated with fixed control inputs over a time interval with known initial states. Given the actuation constraints $v_{max}$ and $\omega_{max}$ of the robot, we uniformly generate $N_{v}\times N_{\omega}$ samples from $[0,v_{max}]\times[-\omega_{max},\omega_{max}]$ as the finite set of control inputs. Subsequently, motion primitives are constructed given the dynamics model, the controls $u$, and time discretization. Collision Checking. We sample a fixed number of points on each motion primitive to conduct collision checks. Since we have an uncertain map, we relax $x\in X_{free}$ to a chance constraint. Let $d_{\kappa}$ be the minimum distance between a test point and the set of Gaussians. Our constraint then amounts to the probability of the distance between the test point and the set of Gaussians in the scene being less than some allowed tolerance $\gamma$ with probability $\eta$ i.e. $P(d_{\kappa}<\gamma)\leq\eta$. When checking for collisions, each sampled point is bounded with a sphere of radius $r_{robot}$ and the radius $r$ of each Gaussian is scaled by a factor of $\lambda_{g}$. The truth value of the test is determined by comparing the distance to all Gaussians with $\gamma:=r_{robot}+\lambda_{g}r$, as illustrated in Fig. 2. Note that setting $\lambda_{g}=3$ is equivalent to the test proposed by [37]. However, unlike the continuous trajectory optimization approach in [37], which requires subsampling to provide soft constraints, our search-based method allows dense checks. We are implicitly assuming that 3DGS accurately captures geometry of the scene and we notice this holds empirically since we initialize Gaussians in the map using the depth measurements from the RGB-D sensor. In the case of collision checking against non-isotropic Gaussians, the principal axis of the Gaussian can be taken as the radius. We develop a GPU-accelerated approach for testing collisions of all sampled points with all Gaussians from the map at once while growing the search tree. This allows the real-time expansion of the search tree.
Tree Search. The cost of each valid motion primitive is defined by Prob. 1. Since the maximum velocity of the robot is bounded by $v_{max}$, we consider the minimum time heuristic as $h(p):=||p_{goal}-p||_{2}/v_{max}$. We use A* to search through the motion primitives tree and keep top $N_{traj}$ candidate trajectories for the information gain maximization.
Information Gain Maximization. A state sequence $\{x_{i}({l\tau/L})\ |\ l\in\{0,\ldots,L\}\}$ containing the end state of each of $L$ segments in the $i$-th trajectory is used for evaluating the information along the trajectory. The information of each state is evaluated according to eq. (9). Finally, the trajectory with the highest information
@@ IV-C Implementation Details
We implemented the proposed method using PyTorch2.4 and the mapping module following [14]. We note that our framework is compatible with other Gaussian splatting approaches if they can meet the compute and latency requirements. The parameters of the mapper were set to the default configuration of [14] for the TUM dataset. For simulation experiments, we used 50 mapping iterations. For real-world experiments, to enable real-time mapping and collision avoidance, we reduced the number of mapping iterations to 10 and pruned Gaussians once per mapping sequence. The mapping and planning modules are both set to 1Hz onboard the robot. For trajectory viewpoint evaluation, we set the threshold for high and low uncertainty Gaussians at half a standard deviation above and below the average magnitude of parameter updates, respectively. We found Gaussians on the ground plane to be particularly noisy in indoor environments due to reflections off the floor. This necessitated the removal of Gaussians on the ground plane for planning. We set the Gaussian region size to 2.5m, $\lambda_{\xi}=1$ and $\lambda_{g}=3$. For planning, we set $N_{v}=3$, $v_{max}=0.6$, $N_{\omega}=5$, $\omega_{max}=0.9$, and time discretization 1s.
@@ V Results
V-A Simulation Experiments & Ablations
Table I: Benchmark experiments on iTHOR test scenes with 100 steps
@@ V-A Simulation Experiments & Ablations
Table II: Ablation experiments on iTHOR test scenes with 100 steps
@@
V-A1 Experiment setup & Baselines
To evaluate our proposed heuristic, we conduct simulation experiments with the AI2-THOR simulator [46] on the iTHOR dataset. The iTHOR dataset consists of indoor scenes including kitchens, bedrooms, bathrooms and living rooms. Each step in the simulator involves a translation and rotation. The agent is allowed to translate 0.5m in both $x$ and $y$ positions. At each allowed position, we sample a set of candidate yaw angles to form a discrete set of possible orientations. These candidate positions, along with their corresponding yaw angles, define the full set of potential viewpoints. Each viewpoint is evaluated using an uncertainty metric, and the one with the highest uncertainty is selected as the next action. We ran experiments on all test scenes from iTHOR with 100 steps each. The simulations were run on a desktop computer with an AMD Ryzen Threadripper PRO 5975WX and NVIDIA RTX A4000 (16GB).
@@ V-A2 Benchmark experiments
The averaged results over all test scenes are presented in Tab. I. Our heuristic performs comparably with FisherRF and outperforms Ensemble across all metrics, while being more than an order of magnitude faster in computation time. Since we are computing the uncertainty directly on the Gaussians instead of the image, we avoid the computation cost of rendering every sampled viewpoint. This efficiency in computation is crucial in a sampling-based planner where we evaluate potentially hundreds of viewpoints in each planning iteration.
@@ V-B Planner Performance
We evaluate the necessity of our proposed GPU-based planning approach in enabling real-time planning. In particular, we conducted experiments on trajectory planning with a 5-meter horizon, evaluating both GPU-accelerated collision checking and a fully CPU-based implementation. As illustrated in Fig. 3, the planning time of CPU-based planner increases significantly with the number of Gaussians. For example, a 40m $\times$ 40m outdoor parking lot can contain around $3\times 10^{6}$ Gaussians. In this case, the CPU-based planner requires 11.65 seconds, whereas the GPU-based planner takes 0.62 seconds, achieving 18$\times$ speedup. The GPU-based collision testing enables parallel growth of the search tree at each layer, along with simultaneous and efficient collision checks between all test points and Gaussians directly on the GPU.
@@
V-C Real-world Experiments & Benchmarks
Table III: Quantitative Results of Real-world Experiments.
Methods Env. Budget (Time) [s] PSNR [dB] $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ RMSE [m] $\downarrow$ mIoUa $\uparrow$
Ensemble Indoor 1 72m2 300 6.27 0.204 0.839 2.051 0.071
FisherRF 16.80 0.665 0.393 0.242 0.314
RTGuIDE 17.83 0.737 0.334 0.202 0.338
@@ V-C Real-world Experiments & Benchmarks
Ensemble Indoor 2 189m2 600 10.41 0.269 0.787 2.410 0.156
FisherRF 14.88 0.588 0.451 1.170 0.248
RTGuIDE 16.40 0.711 0.299 0.689 0.391
[…]
Ensemble Outdoor 1 252m2 600 10.59 0.392 0.637 1.705 0.549
FisherRF 17.58 0.687 0.349 0.519 0.600
RTGuIDE 20.00 0.744 0.296 0.369 0.622
[…]
Ensemble Outdoor 2 480m2 900 16.79 0.538 0.576 1.422 0.159
FisherRF 19.44 0.680 0.439 0.773 0.177
RTGuIDE 20.27 0.750 0.345 0.399 0.174
[…]
@@ V-C1 Experiment setup & Baselines
To evaluate the effectiveness of our proposed framework, we compared it with two baselines onboard the robot. We first implemented FisherRF [36] by constructing both a voxel map and a Gaussian splatting map online. We detect and cluster frontiers from the voxel map and use A* to generate reference paths to frontier clusters. Camera views along each path are evaluated to compute information gain. To generate collision-free and dynamically feasible trajectories, we generate trajectories along waypoints in the path through MPL [45] in the voxel map considering the unicycle model for robot dynamics. We also implemented an Ensemble method and deployed it on the robot, similar to the framework of FisherRF. In the ensemble approach, uncertainty is computed as described in Sec. V-A1.
We deployed our method and the baselines in four different real-world environments on a Clearpath Jackal robot outfitted with an AMD Ryzen 5 3600 and RTX 4000 Ada SFF. In addition, the platform is equipped with an Ouster OS1 LiDAR for state estimation and a ZED 2i stereo camera for mapping. We used [49] to provide odometry and truncated the depth measurements at 5 meters for all methods in all of our experiments. For the baseline methods, we set the voxel resolution to 5cm. To generate the ground truth data for evaluation, we teleoperated the robot to uniformly sample the environment. The exploration budget for all methods is set to $\sim$5x the time needed to fully survey the environment with teleoperation. We rendered 500 novel views according to the ground truth poses and compared them against the corresponding images.
We computed the same image metrics on the rendered RGB and depth images as in the simulation experiments. To verify the usefulness of the generated Gaussian map representation for downstream tasks in robotics, we also performed an evaluation on the task of semantic segmentation. We used Grounded SAM 2 [50] to obtain segmentation masks of the rendered and groundtruth images and computed the mean Intersection over Union (mIoU).
V-C2 Qualitative results
@@
We evaluated the quality of the Gaussian splatting maps that were constructed in real-time onboard the robot by rendering novel views with a set of test poses. A visualization of the onboard constructed map is shown in Fig. 4, and examples of rendered color images are presented in Fig. 5. Constructed in real time onboard the robot, our Gaussian splatting map provides a detailed representation of the environment. Being able to render photorealistic images from the map benefits downstream tasks such as semantic segmentation, which are analyzed in detail in the next section.
V-C3 Quantitative results
As shown in Tab. III, in real-world scenarios when the operational budget is limited, our proposed framework constructs maps of higher quality than the baselines onboard the robot. To validate that our approach generates maps of reasonable quality, we also present maps generated with the ground truth samples. Results on novel view image rendering show that the maps constructed with our approach achieve 0.8-2.4 higher PSNR than FisherRF and 3.5-11.6 higher PSNR than Ensemble. Furthermore, when evaluated on depth image recovery, our method reduces the RMSE of the rendered depth by 16.5%-48% compared to FisherRF, and by 71.4%-90.2% compared to Ensemble. These results reflect that our method outperforms the baselines in achieving good coverage of the environments while also optimizing for map quality during exploration.

### v3

@@ IV-B1 High-level planner
In contrast to traditional mapping representations, the Gaussian map does not encode occluded and free space. Instead of computing (geometric) frontiers, we use the Gaussian uncertainty estimates to identify regions of the map that should be visited next. We evenly partition the a priori known enclosing space of the desired map into cuboidal regions. At time k{k}, for a region $o$, denote the cardinality of the Gaussians in the region by $M_{k}$. We compute the mean uncertainty of that region $\Omega_{k}=1/{M_{k}}\left(\sum_{\mu\in o}\|\mu_{k}-\mu_{{k-1}}\|_{2}\right).$
We formulate a high-level guidance path that allows us to (i) navigate to regions of high uncertainty; and (ii) utilize the known traversable space to plan long-range trajectories without computationally expensive collision checks. At the high level of our planner, we construct a tree by incrementally adding nodes along the traveled path. The tree consists of two types of nodes: odometry nodes and viewpoint nodes. As the robot moves, it verifies the presence of a nearby odometry node; if none exist, a new odometry node is added to the tree and connected to the nearest existing odometry nodes. At each planning iteration, we sample a fixed number of viewpoints around the identified regions. We compute the shortest viewing distance from the Gaussian region centroid to the optical center of the camera, given the camera intrinsics. Each sampled viewpoint is then assigned the utility computed for that region and connected to the closest odometry node in the tree as a viewpoint node. We then use Dijkstra’s algorithm to find the shortest path from the current robot location in the tree to all the viewpoint nodes in the tree. Finally, we compute the cost-benefit of a path using $\Omega/e^{d}$ where $\Omega$ is the estimated information gain and $d$ is the distance to the node [29]. The maximal cost-benefit path is sent to the trajectory planner.
@@ IV-B2 Trajectory planner
[…]
The first term in eq. (8) encourages additional observations of parts of the scene with high uncertainty. The second term penalizes observation of stable Gaussians in explored areas and encourages exploration of unseen parts of the environment. The two terms are weighted by $\lambda_{\xi}$.
@@
Trajectory Candidate Generation. Given an initial robot state $x_{0}\in\mathcal{X}_{free}$, and goal region $\mathcal{X}_{goal}$, find the control inputs $u(\cdot)$ defined on $0,\tau]$ that solve:
$\min_{u(\cdot),\tau}\ \lambda_{t}\tau+\int_{0}^{\tau}u(t)^{T}u(t)dt$ (9)
@@
Motion Primitive Tree Generation. Inspired by [30], we solve problem 1 by performing a tree search on the motion primitives tree. Motion primitives are generated with fixed control inputs over a time interval with known initial states. Given the actuation constraints $v_{max}$ and $\omega_{max}$ of the robot, we uniformly generate $N_{v}\times N_{\omega}$ samples from $0,v_{max}]\times[-\omega_{max},\omega_{max}]$ as the finite set of control inputs. Subsequently, motion primitives are constructed given the dynamics model, the controls $u$, and time discretization. Collision Checking. We sample a fixed number of points on each motion primitive to conduct collision checks. Since we have an uncertain map, we relax $x\in X_{free}$ to a chance constraint. Let $d_{\kappa}$ be the minimum distance between a test point and the set of Gaussians. Our constraint then amounts to the probability of the distance between the test point and the set of Gaussians in the scene being less than some allowed tolerance $\gamma$ with probability $\eta$ i.e. $P(d_{\kappa}<\gamma)\leq\eta$. When checking for collisions, each sampled point is bounded with a sphere of radius $r_{robot}$ and the radius $r$ of each Gaussian is scaled by a factor of $\lambda_{g}$. The truth value of the test is determined by comparing the distance to all Gaussians with $\gamma:=r_{robot}+\lambda_{g}r$, as illustrated in Fig. 2. Note that setting $\lambda_{g}=3$ is equivalent to the test proposed by [15]. However, unlike the continuous trajectory optimization approach in [15], which requires subsampling to provide soft constraints, our search-based method allows dense checks. We are implicitly assuming that 3DGS accurately captures geometry of the scene and we notice this holds empirically since we initialize Gaussians in the map using the depth measurements from the RGB-D sensor. In the case of collision checking against non-isotropic Gaussians, the principal axis of the Gaussian can be taken as the radius. We develop a GPU-accelerated approach for testing collisions of all sampled points with all Gaussians from the map at once while growing the search tree. This allows the real-time expansion of the search tree.
[…]
Information Maximization. A state sequence $\{x_{i}({l\tau/L})\ |\ l\in\{0,\ldots,L\}\}$ containing the end state of each of $L$ segments in the $i$-th trajectory is used for evaluating the information along the trajectory. The information of each state is evaluated according to eq. (8). Finally, the trajectory with the highest information
@@ IV-C Implementation Details
We implemented the proposed method using PyTorch2.4 and the mapping module following [8]. We note that our framework is compatible with other Gaussian splatting approaches if they can meet the compute and latency requirements. The parameters of the mapper were set to the default configuration of [8] for the TUM dataset. For viewpoint selection experiments, we used 50 mapping iterations. For real-world experiments, to enable real-time mapping and collision avoidance, we reduced the number of mapping iterations to 10 and pruned Gaussians once per mapping sequence. The mapping and planning modules are both set to 1Hz onboard the robot. For trajectory viewpoint evaluation, we set the threshold for high and low uncertainty Gaussians at half a standard deviation above and below the average magnitude of parameter updates, respectively. We found Gaussians on the ground plane to be particularly noisy in indoor environments due to reflections off the floor. This necessitated the removal of Gaussians on the ground plane for planning. We set the Gaussian region size to 2.5m, $\lambda_{\xi}=1$ and $\lambda_{g}=3$. For planning, we set $N_{v}=3$, $v_{max}=0.6$, $N_{\omega}=5$, $\omega_{max}=0.9$, and time discretization 1s.
@@ V Results
V-A Information Metric Evaluation & Ablation
V-A1 Experiment setup & Baselines
To evaluate our proposed heuristic on viewpoint selection, we conducted simulation experiments with the AI2-THOR simulator [31] on the iTHOR dataset. In these experiments, time budget is not imposed, and we run experiments on all test scenes with 100 steps each. The iTHOR dataset consists of indoor scenes including kitchens, bedrooms, bathrooms and living rooms. At each step, the agent is allowed to translate 0.5m in both $x$ and $y$ positions and a discrete set of possible orientations. These candidate positions, along with their corresponding yaw angles, define the full set of potential viewpoints. Each viewpoint is evaluated using an uncertainty metric, and the one with the highest uncertainty is selected as the next action. The simulations were run on a desktop computer with an AMD Ryzen Threadripper PRO 5975WX and NVIDIA RTX A4000 (16GB).
Table I: Information metric benchmark experiments on iTHOR
@@ V-A1 Experiment setup & Baselines
Viewpoint selected using information metric from FisherRF.
Table II: Information metric ablation experiments on iTHOR
@@ V-A2 Benchmark experiments
The averaged results over all test scenes are presented in Tab. I. Our heuristic performs comparably with FisherRF and outperforms Ensemble across all metrics, while being more than an order of magnitude faster in computation time. FisherRF renders images at every sampled viewpoint to approximate the information gain with the full map parameters, while we avoid such computational cost by computing the uncertainty directly on the parameters of the Gaussians. This trades off the approximation accuracy with efficiency. This efficiency in computation is crucial for real-time operation where we evaluate potentially hundreds of viewpoints in each planning iteration.
@@ V-B Planner Performance
V-C Closed-loop Simulation Experiments
Table III: Closed-loop Experiments on MP3D in Unity Simulator
Methods PSNR SSIM LPIPS RMSE Coverage
[dB] $\uparrow$ $\uparrow$ $\downarrow$ [m] $\downarrow$ [%] $\uparrow$
Ensemble 10.71 0.24 0.76 2.14 38.93
FisherRF 15.37 0.49 0.59 1.06 66.40
RTGuIDE 16.32 0.56 0.54 0.71 92.36
GT 18.13 0.65 0.47 0.50
Table IV: Quantitative Results of Real-world Experiments.
Methods Env. Budget (Time) [s] PSNR [dB] $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ RMSE [m] $\downarrow$ Coverage [%] $\uparrow$ mIoUa $\uparrow$
Ensemble
Indoor 1
72m2
300 6.27 0.204 0.839 2.051 98.48 0.071
FisherRF 16.80 0.665 0.393 0.242 99.86 0.314
RTGuIDE 17.83 0.737 0.334 0.202 99.95 0.338
GT 20.56 0.805 0.237 0.158 0.420
Ensemble
Indoor 2
189m2
600 10.41 0.269 0.787 2.410 84.10 0.156
FisherRF 14.88 0.588 0.451 1.170 72.80 0.248
RTGuIDE 16.40 0.711 0.299 0.689 100.00 0.391
GT 18.39 0.803 0.243 0.454 0.422
Ensemble
Outdoor 1
252m2
600 10.59 0.392 0.637 1.705 48.30 0.549
FisherRF 17.58 0.687 0.349 0.519 85.81 0.600
RTGuIDE 20.00 0.744 0.296 0.369 100.00 0.622
GT 22.95 0.828 0.241 0.195 0.631
Ensemble
Outdoor 2
480m2
900 16.79 0.538 0.576 1.422 49.59 0.159
FisherRF 19.44 0.680 0.439 0.773 60.95 0.177
RTGuIDE 20.27 0.750 0.345 0.399 98.35 0.174
GT 24.19 0.835 0.285 0.227 0.246
@@ V-C1 Experiment setup & Baselines
We conducted experiments in a Unity-based simulator with ROS and a simulated Clearpath Jackal robot with an RGBD sensor and ground truth odometry. This set of experiments was designed to evaluate all methods with real-time and closed-loop operation, considering robot dynamics and collision avoidance.
The original implementation of FisherRF [24] does not model dynamics or include real-world experimental results, so we implement a closed-loop version of it. We first construct both a voxel map and a Gaussian splatting map online. We detect and cluster frontiers from the voxel map and use A* to generate reference paths to frontier clusters. Camera views along each path are evaluated to compute information gain. To generate collision-free and dynamically-feasible trajectories, we generate trajectories along waypoints in the path through MPL [30] in the voxel map. We similarly implement an Ensemble method as described in Sec. V-A1.
We obtained the ground truth (GT) by tele-operating the robot to uniformly survey the environment. The collected data was used to build a 3DGS map for GT and also served as test poses for the other methods. This represents the upper bound of the evaluation metrics with the Gaussian splatting method. The exploration budget for all methods is set to approximately 5 times the duration needed to fully survey the environment with tele-operation. For the baseline methods, we set the voxel resolution to 5cm. We rendered 500 novel views at the ground truth test poses for evaluation. The same image metrics on the rendered RGB and depth images are used as in Sec. V-A. We evaluated all methods on 5 MP3D [34] scenes [17DRP5sb8fy, HxpKQynjfin, 2t7WUuJeko7, 8194nk5LbLH, YVUC4YcDtcY] with a time budget of 1500s, 480s, 1500s, 1200s and 1800s respectively. These scenes consist of multi-room indoor environments with varying size and clutter.
V-C2 Results
The averaged statistics of the experiments are presented in Table III. We note that there are signifcant errors in the meshes in MP3D which affect the reconstruction results for all methods including the ground truth. Our proposed method achieves 0.95dB and 5.61dB higher PSNR than FisherRF and Ensemble. The significantly higher coverage demonstrates the importance of efficient viewpoint selection in real-time missions with limited operational budget.
V-D Real-world Experiments & Benchmarks
V-D1 Experiment setup & Baselines
To evaluate the effectiveness of our proposed framework in real-world settings, we compared it with two baselines onboard the robot. We followed the same setup as in the closed-loop simulation experiments for the baselines and the ground truth GT.
We deployed our method and the baselines in four different real-world environments on a Clearpath Jackal robot outfitted with an AMD Ryzen 5 3600 and RTX 4000 Ada SFF. In addition, the platform is equipped with an Ouster OS1 LiDAR for state estimation and a ZED 2i stereo camera for mapping. We used [35] to provide odometry and truncated the depth measurements at 5 meters for all methods in all of our experiments. The same evaluation setup from Sec. V-C is employed. To verify the usefulness of the generated Gaussian map representation for downstream tasks in robotics, we also performed an evaluation on the task of semantic segmentation. We used Grounded SAM 2 [36] to obtain segmentation masks of the rendered and groundtruth images and computed the mean Intersection over Union (mIoU).
V-D2 Qualitative results
@@
We evaluated the quality of the Gaussian splatting maps that were constructed in real-time onboard the robot by rendering novel views with a set of test poses. A visualization of the onboard constructed map is shown in Fig. 4, and examples of rendered color images are presented in Fig. 5. Constructed in real time onboard the robot, our Gaussian splatting map provides a detailed representation of the environment. Being able to render photorealistic images from the map benefits downstream tasks such as semantic segmentation, which are analyzed in the next section.
V-D3 Quantitative results
As shown in Tab. IV, in real-world scenarios when the operational budget is limited, our proposed framework constructs maps of higher quality than the baselines onboard the robot. To validate that our approach generates maps of reasonable quality, we also present maps generated with the ground truth samples. Results on novel view image rendering show that the maps constructed with our approach achieve 0.8-2.4 higher PSNR than FisherRF and 3.5-11.6 higher PSNR than Ensemble. Furthermore, when evaluated on depth reconstruction, our method reduces the RMSE of the rendered depth by 16.5%-48% compared to FisherRF, and by 71.4%-90.2% compared to Ensemble. These results reflect that our method outperforms the baselines in achieving good coverage of the environments while also optimizing for map quality during exploration.